Shift report 20000901 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Fri Sep 01 2000 - 16:01:38 EDT

  • Next message: http account: "Shift report 20000902 16:00-24:00"

    ------------------------------------------------------------
    Date: Fri Sep  1 16:01:35 2000, Shift: 08:00-16:00
    Supervisor: ZM, with: EJ
    Log book#5 pages 71 - 74
    
    Shift Summary:
    ------------------------------------------------------------
    1. Continous beam until 11:34 - parasitic studies,
       13:30 - ring access. At 15:00 power outage in 1000P.
       16:00 recent info - transformer problem. 
    2. Run#2494 collected 234K trig.#1
       Run#2495 collected 166K trig.#1
       YB&KH set BBdicr. at 100mV (all small tubes and big R31)
       other big tubes - 150 mV, Expert says that BFS platform
       canot moved. 
       RD check the magnets poles: D5 points North whe polarity is 
       set to A, D2 points South when polarity is set B. Connection
        of D5 to the Keithley readout was fixed. 
    3. No detector problems.
       
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Fri Sep 01 2000 - 16:01:59 EDT